I need to combine these mixed fractions. 6 2/9 minus 3 5/12. I keep getting 6 23/36 and everything is right but the 23 should be 29.
Answer by stanbon(75887)

I need to combine these mixed fractions. 6 2/9 minus 3 5/12.
----
6 2/9 = (9*6+2)/9 = 56/9
------
3 5/12 = (12*3+5)/12 = 41/12
------
lcd = 36
----
(56/9)(4/4) = 224/36
---
(41/12)(3/3) = 123/36
-----
(224/36)-(123/36) = 101/36 = 2 29/36
======================================
Cheers,
Stan H.